OpenStudy (anonymous):

A line k passes through the points (4,2) and (-4,4). A second line h passes through the points (-8,1) and (8, -3). Is line k parallel to line h? Why or why not?

OpenStudy (anonymous):

slope of line k = 2/-8 = -1/4 slope of line h = -4/16 = -1/4 both slope are equal , so both lines are parallel